OpenStudy (anonymous):

a ball is thrown vertically upward from the ground with an initial speed of 25 m/s; at the same instant another ball is dropped from a building 15 m high. after how long will the balls be at the same height? what are their speeds when they are at the same height?

OpenStudy (lgbasallote):

object 1 \[\Large X_{1_F} - X_{1_i} = V_{1_i}t+ \frac 12 a_{1}t^2\] \[\Large \implies X_{1_F} = V_{1_i}t + X_{1_i}+ \frac 12 a_{1}t^2\] object 2 \[\Large \implies X_{2_F} - X_{2_i} = V_{2_i}t + \frac 12 a_{2}t^2\] \[\Large \implies X_{2_F} = V_{2_i}t + - X_{2_i} +\frac 12 a_{2}t^2\] where X is the distance i - initial f = final t = time V = velocity do you agree with that?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

an ideal way is to draw a diagram first..

OpenStudy (lgbasallote):

so when they meet then the X_1F and X_2F would be equal

OpenStudy (lgbasallote):

since final distance of object 1 = final distance of object 2

OpenStudy (lgbasallote):

\[\large V_{1_i} t + X_{1_i} + \frac 12 a_1 t^2 = V_{2_i} t + X_{2_i} + \frac 12 a_2 t^2\] right?

mathslover (mathslover):

right good omni